# Reconsidering the Performance of GAE in Link Prediction

### v3

@@ Abstract.
Recent advancements in graph neural networks (GNNs) for link prediction have introduced sophisticated training techniques and model architectures. However, reliance on outdated baselines may exaggerate the benefits of these new approaches. To tackle this issue, we systematically explore Graph Autoencoders (GAEs) by applying model-agnostic tricks in recent methods and tuning hyperparameters. We find that a well-tuned GAE can match the performance of recent sophisticated models while offering superior computational efficiency on widely-used link prediction benchmarks. Our approach delivers substantial performance gains on datasets where structural information dominates and feature data is limited. Specifically, our GAE achieves a state-of-the-art Hits@100 score of 78.41% on the ogbl-ppa dataset. Furthermore, we examine the impact of various tricks to uncover the reasons behind our success and to guide the design of future methods. Our study emphasizes the critical need to update baselines for a more accurate assessment of progress in GNNs for link prediction. Our code is available at https://github.com/GraphPKU/Refined-GAE. [^1] This paper has been accepted for presentation at the 34th ACM International Conference on Information and Knowledge Management (CIKM 2025).
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Neural networksccs: Mathematics of computing Graph algorithms
@@ 1. Introduction
Link prediction is a fundamental problem in graph learning with applications in recommendation systems (Zhang and Chen, 2020), drug discovery (Souri et al., 2022), and knowledge graphs (Zhu et al., 2021). GNNs have achieved strong performance in these domains and are widely adopted for link prediction tasks. Among GNNs, GAE (Kipf and Welling, 2016) is a well-known model that predicts link probabilities by computing inner products of node representations learned through a Message Passing Neural Network (MPNN).
@@
However, Zhang and Chen (2018) found that GAE’s expressiveness is limited. For example, as illustrated in Figure 1a, GAE produces identical predictions for links $(v_{1},v_{2})$ and $(v_{1},v_{3})$, despite their differing structural contexts. This limitation arises because GAE computes node representations independently, disregarding the structural relationships between nodes. To address this challenge, various methods have been proposed to enhance expressiveness and better capture graph structure, such as labeling tricks (Zhang and Chen, 2018), paths between nodes (Zhu et al., 2021), and neighborhood overlaps (Liang et al., 2024; Yun et al., 2021; Chamberlain et al., 2023; Wang et al., 2024b). These approaches often introduce complex designs with increased time and space complexity, claiming superior performance. Although expressiveness improvement caused by architecture changes is important, we observe that holistic optimization of existing GAE components and meticulous hyperparameter tuning also contribute significantly to these models’ success.
@@
This observation motivates a fair re-evaluation: baseline GAEs must also be applied with such optimization to be compared with extensively tuned models. We therefore reimplemented GAE, systematically applying principled, relatively simple enhancements and meticulous hyperparameter tuning, practices common in modern GNN development. Our approach also introduces a flexible input strategy and key architectural optimizations. These empirical efforts are supported by theoretical analysis; this analysis of GAE with linear propagation and unique inputs reveals an inherent capacity for processing pairwise neighborhood information, akin to capturing generalized common neighbor information. The theoretical grounding, further reinforced by empirical findings such as trained learnable embeddings tending to remain near-orthogonal, validates the effectiveness achieved through optimized GAE.
@@
Through extensive experiments on Planetoid (Sen et al., 2008) and Open Graph Benchmark (Hu et al., 2020) datasets, we demonstrate that properly Optimized GAE, with its core simplicity largely intact, can match or surpass SOTA models. As detailed in Table 2, our Optimized GAE achieves top-two performance on all listed datasets and attains the leading average rank compared against all baselines, notably yielding an average improvement of 6.4% over the strongest NCN baseline (Wang et al., 2024b). Notably, it establishes a new SOTA Hits@100 score of 78.41% on the large-scale ogbl-ppa dataset. These results demonstrate that an optimized GAE model can serve as a powerful and more efficient alternative, comparable to complex GNN architectures that explicitly inject pairwise information. The general applicability of these optimization principles is confirmed by similar performance improvements when applied to NCN (Wang et al., 2024b). This highlights the benefits of a carefully refined base model. Finally, comprehensive ablation studies validate the contribution of each design choice of GAE, providing valuable observations to guide the future development of link prediction models.
@@
  • We demonstrate that a GAE model with systematic optimization can achieve link prediction performance competitive with, or even better than, SOTA models. This finding shows that systematic optimization of simpler architectures can yield strong results, supporting the critical re-evaluation of basic models to establish more effective and efficient baselines in the field.
  • We provide a comprehensive analysis of the crucial design choices and optimization strategies spanning input representation, architectural configurations, and hyperparameter tuning. We detail their effects to offer actionable guidelines for the development of efficient and powerful link prediction models.
@@ 2. Related Works
The GAE (Kipf and Welling, 2016), an early GNN-based method, uses an MPNN to generate node embeddings and predicts links using inner product similarity. However, its expressiveness is limited because it does not explicitly model pairwise relationships. To address this, SEAL (Zhang and Chen, 2018) improves expressiveness by extracting and processing local subgraphs around target links with a GNN. Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023) process the entire graph, using higher-order common neighbor information for improved global modeling. NCN (Wang et al., 2024b) further develops pairwise modeling by learning embeddings for common neighbors instead of using manually defined features. Other research has focused on generalizing heuristic methods. NBFNet (Zhu et al., 2021) unifies path-based heuristics using the Bellman-Ford algorithm, while Mao et al. (2023) categorizes heuristics into local and global ranges for broader structural modeling. LPFormer (Shomer et al., 2024) uses an adaptive transformer to dynamically select relevant heuristics for each dataset. Similarly, Link-MOE (Ma et al., [n. d.]) uses an ensemble of link predictors to flexibly combine multiple heuristics.
Research has also addressed data-related challenges affecting link prediction. For example, LTLP (Wang et al., 2024a) reduces bias from long-tailed distributions to improve performance on imbalanced datasets. Zhu et al. (2024) studied the impact of homophily and heterophily.
Baseline Evaluation and Optimization in GNNs. As new GNN models are introduced, recent studies suggest caution with evaluation practices that might exaggerate the benefits of complex models. For link prediction, Li et al. (Li et al., 2024) identified key problems such as poorly tuned baselines and unrealistic negative sampling in evaluations, which can hide the true performance differences between models. Similar concerns exist in other graph tasks. Luo et al. (2024) showed that classic GNNs with carefully tuned under standard methods can perform even better than complex Graph Transformers. This applies to node-level (Luo et al., 2024) and graph-level tasks (Luo et al., 2025). These findings question the assumption of the necessity of complex architectures.
@@ 3. Preliminaries
GAE. GAEs (Kipf and Welling, 2016) are a type of model used for link prediction. GAEs learn low-dimensional node embeddings that preserve graph structure and node attributes. These embeddings are then used to predict link likelihoods. A GAE has two main components:
@@ 4.1. GAE’s Inherent Pairwise Structure Modeling
However, GAE applications for link prediction typically use non-permutation-invariant inputs, departing from the assumption of uniform initial labeling. This key difference arises in two main ways: nodes may have unique features, or unique initial embeddings (learnable or distinctly initialized) are used. Either approach gives each node a unique starting signature. This fundamentally changes the GAE’s operating conditions compared to those assumed in analyses that rely on initial permutation invariance. Our theoretical exploration of GAE’s pairwise processing ability starts from this practical consideration.
@@
In conclusion, using unique initial node features or embeddings allows GAEs to inherently perform complex pairwise information processing. They can therefore effectively capture common neighbor signals and assess node environment compatibility, offering more detailed insights for link prediction. This perspective complements traditional analyses of expressiveness that focus on purely structural information.
@@ 4.2.2. Key Architectural Modules of Graph Autoencoders
Non-Linear Activation (and its strategic absence). Section 4.1 highlighted that inner products of MPNN-derived embeddings can capture pairwise relationships. However, standard GNN designs often apply non-linear activations after convolution operation, resulting in updates like $H^{(l+1)}=\sigma(\text{Conv}_{l+1}(H^{(l)},A))$. Such intermediate non-linearities can disrupt the direct structural information encoded in dot products. Therefore, we strategically remove non-linear activations between sequential MPNN layers. Each layer therefore performs a linear transformation. This approach is similar in principle to linear GNNs like SGC (Wu et al., 2019), where the update rule is $Z^{(l+1)}=\text{Conv}_{l+1}(Z^{(l)},A)$. Stacking these linear layers allows the final encoder output to keep structural properties through inner products. The non-linearity is then introduced in the decoder stage. Several other recent works have also effectively used linear propagation for link prediction, including LightGCN (He et al., 2020), HL-GNN (Zhang et al., 2024), YinYanGNN (Wang et al., 2025), and MPLP (Dong et al., 2023).
After the encoder produces node embeddings, the decoder can use the dot product followed by a sigmoid function to predict edge likelihoods. A more expressive alternative is to use an MLP, which can learn non-linear interactions between node embeddings. This is especially important if the encoder’s MPNN layers are linear, as the MLP in the decoder can then introduce the needed non-linear transformations to model complex relationships. Similar to their use in the encoder’s MPNN layers, initial residual connections also help in training deeper MLPs in the decoder. They help ensure stable training and effective information flow through the predictor’s layers.
4.2.3. Revisiting Standard Hyperparameter Settings
Network Depth. The depth of the network plays an important role in determining its ability to capture complex patterns. However, GNNs have challenges with increasing depth, especially oversmoothing (Li et al., 2018). To address the challenge, we adopt the following approach: our strategy uses shallow MPNN encoder layers yet deep MLP decoder layers. This design allows the model to retain high expressive power because the MLP learns complex relational patterns independently.
5. Experiments
5.1. Experiment Setup
@@
Optimized-GAE $\underline{88.17\pm 0.93}$ $\underline{92.40\pm 1.23}$ 80.09 $\pm$ 1.72 $\underline{66.11\pm 0.35}$ 78.41 $\pm$ 0.83 $\underline{88.74\pm 0.06}$ 94.43 $\pm$ 0.57
Table 3. Side-by-side comparison of reported performance of NCN and performance after applying our optimization techniques(denoted with *).
@@
Datasets We evaluate our model on three datasets from Planetoid (Sen et al., 2008): Cora, Citeseer, and Pubmed; and four datasets from OGB (Hu et al., 2020): ogbl-collab, ogbl-ddi, ogbl-ppa, and ogbl-citation2.
Baselines. We use the official OGB results for GAE as baselines. For a broader comparison, we also include heuristic methods such as Common Neighbors (CN) (Barabási and Albert, 1999), Adamic-Adar (AA) (Adamic and Adar, 2003), and Resource Allocation (RA) (Zhou et al., 2009), as well as GNN models like SEAL (Zhang and Chen, 2018), NeoGNN (Yun et al., 2021), NBFNet (Zhu et al., 2021), BUDDY (Chamberlain et al., 2023), NCN (Wang et al., 2024b) and MPLP+ (Dong et al., 2023). Baseline results are taken from three established studies (Wang et al., 2024b; Dong et al., 2023; Hu et al., 2020).
Experiment settings. For the Planetoid datasets, we use a 70%-10%-20% train-validation-test split and evaluate performance using Hits@100 following prior works. For OGB datasets, we use the official data splits and evaluation metrics. Guided by the input representation strategy, our main Optimized GAE model incorporates learnable node embeddings for the ogbl-ddi, ogbl-ppa, and ogbl-citation2 datasets. Further experimental details are in Appendix A. Noted on ogbl-collab, different evaluation strategies have been used in previous work. Some studies (Zhang and Chen, 2018; Yun et al., 2021; Hu et al., 2020; Chamberlain et al., 2023; Wang et al., 2024b; Dong et al., 2023) include validation edges only during test time. Others (Wang et al., 2025; Zhang et al., 2024; Wang et al., 2021) add the validation set to the training data, following current OGB guidelines. For our main results, we use the first strategy to align with the most common approach in the literature.
5.2. Main Results
The main results are presented in Table 2. Our model achieves SOTA on 3 datasets and top-2 on 4 datasets. The consistently strong performance reflects an average improvement of 91.5% over the original GAE baseline. For instance, on the structure-reliant ogbl-ppa and ogbl-ddi datasets, where our input representation technique is particularly effective, performance improves from 18.67% to 78.41% and from 37.07% to 94.43%. Furthermore, Optimized-GAE shows an average improvement of 6.4% over NCN (Wang et al., 2024b), a competitive SOTA model, surpassing it on six out of seven datasets. Notably, on ogbl-ppa, our model’s Hits@100 score of 78.41% not only exceeds the strong MPLP+ baseline by a significant 20.2% margin but also outperforms the OGB leaderboard leader by 2.4%. These results support our main claim: a properly optimized GAE can match or exceed the performance of current models that use more complex methods, validating GAE’s potential in current link prediction tasks.
5.3. Ablation Study

### v4

@@ Abstract.
Recent advancements in graph neural networks (GNNs) for link prediction have introduced sophisticated training techniques and model architectures. However, reliance on outdated baselines may exaggerate the benefits of these new approaches. To tackle this issue, we systematically explore Graph Autoencoders (GAEs) by applying model-agnostic tricks in recent methods and tuning hyperparameters. We find that a well-tuned GAE can match the performance of recent sophisticated models while offering superior computational efficiency on widely used link prediction benchmarks. Our approach delivers substantial performance gains on datasets where structural information dominates and feature data is limited. Specifically, our GAE achieves a state-of-the-art (SOTA) Hits@100 score of 78.41% on the ogbl-ppa dataset. Furthermore, we examine the impact of various tricks to uncover the reasons behind our success and to guide the design of future methods. Our study emphasizes the critical need to update baselines for a more accurate assessment of progress in GNNs for link prediction. Our code is available at https://github.com/GraphPKU/Refined-GAE.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761108isbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Neural networksccs: Mathematics of computing Graph algorithms
@@ 1. Introduction
Link prediction is a fundamental problem in graph learning with applications in recommendation systems (Zhang and Chen, 2020), drug discovery (Souri et al., 2022), and knowledge graphs (Zhu et al., 2021). GNNs have achieved strong performance in these domains and are widely adopted for link prediction tasks. Among GNNs, GAE (Kipf and Welling, 2016a) is a well-known model that predicts link probabilities by computing inner products of node representations learned through a Message Passing Neural Network (MPNN).
@@
However, Zhang and Chen (2018) found that GAE’s expressiveness is limited. For example, as illustrated in Figure 1a, GAE produces identical predictions for links $(v_{1},v_{2})$ and $(v_{1},v_{3})$, despite their differing structural contexts. This limitation arises because GAE computes node representations independently, disregarding the structural relationships between nodes. To address this challenge, various methods have been proposed to enhance expressiveness and better capture graph structure, such as labeling tricks (Zhang and Chen, 2018), paths between nodes (Zhu et al., 2021), and neighborhood overlaps (Liang et al., 2024; Yun et al., 2021; Chamberlain et al., 2023; Wang et al., 2024b). These approaches often introduce complex designs with increased time and space complexity, claiming superior performance. Although the expressiveness improvement caused by architecture changes is important, we observe that holistic optimization of existing GAE components and meticulous hyperparameter tuning also contribute significantly to these models’ success.
@@
This observation motivates a re-evaluation of GAE that moves beyond simple baseline comparisons to instead explore the true potential unlocked when this foundational architecture is enhanced with modern optimization. We therefore reimplemented GAE, systematically applying principled, relatively simple enhancements and meticulous hyperparameter tuning, practices common in modern GNN. Our approach also introduces a flexible input strategy and key architectural optimizations. These empirical efforts are supported by theoretical analysis; this analysis of GAE with linear propagation and unique inputs reveals an inherent capacity for processing pairwise neighborhood information, akin to capturing generalized common neighbor information. The theoretical grounding, further reinforced by empirical findings such as trained learnable embeddings tending to remain near-orthogonal, validates the effectiveness achieved through optimizing GAE.
@@
Through extensive experiments on Planetoid (Sen et al., 2008) and Open Graph Benchmark (Hu et al., 2020) datasets, we demonstrate that properly Optimized GAE, with its core simplicity largely intact, can match or surpass SOTA models. As detailed in Table 3, our Optimized GAE achieves top-two performance on all listed datasets and attains the leading average rank compared against all baselines, notably yielding an average improvement of 6.4% over the strongest NCN baseline (Wang et al., 2024b). Notably, it establishes a new SOTA Hits@100 score of 78.41% on the large-scale ogbl-ppa dataset. These results demonstrate that an optimized GAE model can serve as a powerful and more efficient alternative, comparable to complex GNN architectures that explicitly inject pairwise information. The general applicability of these optimization principles is confirmed by similar performance improvements when applied to NCN (Wang et al., 2024b), highlighting the benefits of a carefully refined base model. Finally, comprehensive ablation studies validate the contribution of each design choice of GAE, providing valuable observations to guide the future development of link prediction models.
@@
  • We affirm the value of re-evaluating foundational baselines by demonstrating that a GAE, enhanced through principled optimizations, achieves SOTA performance that surpasses many complex models. This result provides a counterpoint to the trend of ever-increasing model complexity, establishing that significant gains can be unlocked by refining simple architectures.
  • We distill our comprehensive analysis into a set of principles for building powerful and efficient GAE-based link prediction models. These guidelines, covering key choices from input representation to architectural configuration, offer an actionable roadmap for the future development.
@@ 2. Related Works
The GAE (Kipf and Welling, 2016a), an early GNN-based method, uses an MPNN to generate node embeddings and predicts links using inner product similarity. However, its expressiveness is limited because it does not explicitly model pairwise relationships. To address this, SEAL (Zhang and Chen, 2018) improves expressiveness by extracting and processing local subgraphs around target links with a GNN. Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023) process the entire graph, using higher-order common neighbor information for improved global modeling. NCN (Wang et al., 2024b) further develops pairwise modeling by learning embeddings for common neighbors instead of using manually defined features. Other research has focused on generalizing heuristic methods. NBFNet (Zhu et al., 2021) unifies path-based heuristics using the Bellman-Ford algorithm, while Mao et al. (2023) categorizes heuristics into local and global ranges for broader structural modeling. LPFormer (Shomer et al., 2024) uses an adaptive transformer to dynamically select relevant heuristics for each dataset. Similarly, Link-MOE (Ma et al., [n. d.]) uses an ensemble of link predictors to flexibly combine multiple heuristics. Another line of research elevates link prediction to a probabilistic generative task, a departure from the methods that explicitly model pairwise relations. Starting with the Variational Graph Autoencoder (VGAE) (Kipf and Welling, 2016b), subsequent work has refined this framework, employing methods like incorporating multi-scale information (Guo et al., 2022) or improving generalizability via graph masking (Tan et al., 2023).
Research has also addressed data-related challenges affecting link prediction. For example, LTLP (Wang et al., 2024a) reduces bias from long-tailed distributions to improve performance on imbalanced datasets. Zhu et al. (2024) studied the impact of homophily and heterophily. While our study focuses on the static, homogeneous setting, we note that parallel research addresses the challenges in dynamic graphs, such as TGAT (Xu et al., 2020), and in heterogeneous graphs, such as RGCN (Schlichtkrull et al., 2018).
Baseline Evaluation and Optimization in GNNs. As new GNN models are introduced, recent studies suggest caution with evaluation practices that might exaggerate the benefits of complex models. For link prediction, Li et al. (Li et al., 2024) identified key problems such as poorly tuned baselines and unrealistic negative sampling in evaluations, which can hide true performance differences between models. Similar concerns exist in other graph tasks. Luo et al. (2024) showed that classic GNNs with carefully tuned under standard methods can perform better than complex Graph Transformers. This applies to node-level (Luo et al., 2024) and graph-level tasks (Luo et al., 2025), questioning the necessity of complex architectures. Complementary to their findings, our work offers a deep dive into link prediction, uncovering specific design choices for GAE beyond general hyperparameter tuning.
@@ 3. Preliminaries
GAE. GAEs (Kipf and Welling, 2016a) are a type of model used for link prediction. GAEs learn low-dimensional node embeddings that preserve graph structure and node attributes. These embeddings are then used to predict link likelihoods. A GAE has two main components:
@@ 4.1. GAE’s Inherent Pairwise Structure Modeling
However, GAE applications for link prediction typically use non-permutation-invariant inputs, departing from the assumption of uniform initial labeling. This key difference arises in two main ways: nodes may have unique features, or unique initial embeddings (learnable or distinctly initialized). Either approach gives each node a unique starting signature. This fundamentally changes the GAE’s operating conditions compared to those assumed in analyses that rely on initial permutation invariance. Our theoretical exploration starts from this practical consideration.
@@
In conclusion, using unique initial node features or embeddings allows GAEs to inherently perform pairwise information processing. They can therefore effectively capture common neighbor signals and assess node environment compatibility. This perspective complements traditional analyses of expressiveness that focus on purely structural information.
@@ 4.2.2. Key Architectural Modules of Graph Autoencoders
MPNN Layers $\{2,\,3,\,4\}$ 4 4 4 2 4 2 3
Hidden Dimension $\{256,\,512,\,1024\}$ 1024 1024 512 1024 512 512 256
Batch Size [2048, 65536][2048,\,65536] 2048 4096 4096 8192 16384 65536 65536
Dropout $\{0.2,0.4,0.6\}$ 0.6 0.6 0.4 0.6 0.2 0.2 0.2
MLP Layers $\{2,4,5,8\}$ 4 2 2 8 5 5 5
@@
Non-Linear Activation (and its strategic absence). Section 4.1 highlighted that inner products of MPNN-derived embeddings can capture pairwise relationships. However, standard GNN designs often apply non-linear activations after the convolution operation, resulting in updates like $H^{(l+1)}=\sigma(\text{Conv}_{l+1}(H^{(l)},A))$. Such intermediate non-linearities can disrupt the direct structural information encoded in dot products. Therefore, we strategically remove non-linear activations between sequential MPNN layers. Each layer therefore performs a linear transformation. This approach is similar in principle to linear GNNs like SGC (Wu et al., 2019), where the update rule is $Z^{(l+1)}=\text{Conv}_{l+1}(Z^{(l)},A)$. Stacking these linear layers allows the final encoder output to keep structural properties through inner products. The non-linearity is then introduced in the decoder stage. Several other recent works have also effectively used linear propagation for link prediction, including LightGCN (He et al., 2020), HL-GNN (Zhang et al., 2024), YinYanGNN (Wang et al., 2025), and MPLP (Dong et al., 2023).
After the encoder produces node embeddings, the decoder can use the dot product followed by a sigmoid function to predict edge likelihoods. A more expressive alternative is to use an MLP, which can learn non-linear interactions between node embeddings. This is especially important if the encoder’s MPNN layers are linear, as the MLP can then introduce non-linear transformations to model complex relationships. Similar to their use in the encoder’s MPNN layers, initial residual connections also help in training deeper MLPs in the decoder. They help ensure stable training and effective information flow through the predictor’s layers.
@@
Optimized GAE $\underline{88.17\pm 0.93}$ $\underline{92.40\pm 1.23}$ 80.09 $\pm$ 1.72 $\underline{66.11\pm 0.35}$ 78.41 $\pm$ 0.83 $\underline{88.74\pm 0.06}$ 94.43 $\pm$ 0.57
Table 4. Comparison of reported performance of NCN and performance after applying our optimization (denoted with *).
@@
4.2.3. Revisiting Standard Hyperparameter Settings
Network Depth. The depth of the network plays an important role in determining its ability to capture complex patterns. However, GNNs have challenges with increasing depth, especially oversmoothing (Li et al., 2018). To address the challenge, we adopt the following approach: we use shallow MPNN encoder layers yet deep MLP decoder layers. This design allows the model to retain high expressive power because the MLP learns complex patterns independently.
5. Experiments
5.1. Experiment Setup
5.1.1. Datasets and Baselines
We evaluate our model on three datasets from Planetoid (Sen et al., 2008): Cora, Citeseer, and Pubmed; and four datasets from OGB (Hu et al., 2020): ogbl-collab, ogbl-ddi, ogbl-ppa, and ogbl-citation2.
As baselines, we use the official OGB results for GAE. For a broader comparison, we also include heuristic methods such as Common Neighbors (CN) (Barabási and Albert, 1999), Adamic-Adar (AA) (Adamic and Adar, 2003), and Resource Allocation (RA) (Zhou et al., 2009), as well as GNN models like SEAL (Zhang and Chen, 2018), NeoGNN (Yun et al., 2021), NBFNet (Zhu et al., 2021), BUDDY (Chamberlain et al., 2023), NCN (Wang et al., 2024b) and MPLP+ (Dong et al., 2023). Baseline results are taken from three established studies (Wang et al., 2024b; Dong et al., 2023; Hu et al., 2020).
5.1.2. Implementation Details
For the Planetoid datasets, we use a 70%-10%-20% train-validation-test split, while for OGB datasets, we use the official data splits and evaluation metrics. Our main Optimized GAE model incorporates learnable node embeddings for the ogbl-ddi, ogbl-ppa, and ogbl-citation2 datasets. For ogbl-collab, we follow the common evaluation strategy of including validation edges only during test time (Zhang and Chen, 2018; Yun et al., 2021; Hu et al., 2020; Chamberlain et al., 2023; Wang et al., 2024b; Dong et al., 2023).
All experiments are conducted on Nvidia 4090 GPUs, with Nvidia A800 GPUs used for the ogbl-citation2 dataset due to its higher computational requirements. We train for up to 500 epochs using a 1:3 negative sampling ratio, and all results are averaged over 5 runs with different random seeds. Hyperparameters were tuned using Bayesian optimization to maximize validation performance, with the final settings for each dataset summarized in Table 2.